OpenStudy (anonymous):

What is the point-slope form of the equation of a line that passes through the point (1, 4) and has a slope of −3? A. y-1=-3(x-4) B. y+4=3(x=1) C. y+4=-3(x+1) D. y-4=-3(x-1)

OpenStudy (jojo4eva):

D.

OpenStudy (jojo4eva):

point slope form is in the form of y-y1=m(x-x1) This would turn into y-4=-3(x-1) which gives you answer choice D.

OpenStudy (jojo4eva):

Understood??

OpenStudy (anonymous):

yup thanks

OpenStudy (jojo4eva):

No problem if you have any other questions just tag me